# Revisiting XL-MIMO Channel Estimation: When Dual-Wideband Effects Meet Near Field

### v2

@@ IV-B Column-Wise Hierarchical Sparse Prior Model
To capture the global and local common-delay sparsity characteristics of $\mathbf{X}$, the column-wise hierarchical prior model is characterized as
$p(\mathbf{X},\boldsymbol{\Gamma},\boldsymbol{\Omega},\mathbf{S})=p(\mathbf{X}|\boldsymbol{\Gamma})p(\boldsymbol{\Gamma}|\boldsymbol{\Omega},\mathbf{S})p(\boldsymbol{\Omega})p(\mathbf{S}),$ (22)
where the first layer is modeled as a complex Gaussian distribution with mean zero and variance $[\boldsymbol{\Gamma}]_{n,k}=\gamma_{n,k}^{-1}$, i.e.,
$p(\mathbf{X}|\boldsymbol{\Gamma})=\prod_{n=1}^{N}\prod_{k=1}^{K}\mathcal{CN}(x_{n,k};0,\gamma_{n,k}^{-1}).$ (23)
In the second layer, each $\gamma_{n,k}$ is assigned a conditionally Bernoulli distribution given by
@@
$p(\boldsymbol{\Omega})=\prod_{i=1}^{2}\prod_{k=1}^{K}p(\alpha_{k}^{i})=\prod_{i=1}^{2}\prod_{k=1}^{K}\mathrm{Ga}(\alpha_{k}^{i};a_{i},b_{i}),$ (25)
In the third layer, to capture the global block sparsity, the variance state variable $\mathbf{S}$ is modeled as a MRF
@@
where $\mathcal{D}_{n,k}$ denotes the element set of neighbor nodes of $s_{n,k}$ with $u(s_{n,k},s_{n^{\prime},k^{\prime}})=\exp(\varpi s_{n,k}s_{n^{\prime},k^{\prime}})$, and $v(s_{n,k})=\exp(-\eta s_{n,k})$; $\varpi$ and $\eta$ are the parameters of $p(\mathbf{S})$.
In the following, we discuss how prior models in (22)-(26) capture the structured sparsity in $\mathbf{X}$. 1) Variance-Driven Sparsity: The variance $\gamma_{n,k}^{-1}$ controls the sparsity of $x_{n,k}$, with values approaching zero effectively setting $x_{n,k}$ to zero. 2) Global Block Sparsity: The MRF models the variance states, encouraging block patterns while suppressing isolated coefficients that deviate from their neighbors. 3) Local Common-Delay Sparsity: As described in (24) and (25), the common-delay sparsity mechanism is introduced to ensure that, for a fixed $k$, all $x_{n,k}$ sharing the same variance state are governed by a common variance precision parameter. This enforces a unified sparsity structure among components associated with the same delay.
IV-C Bayesian Inference
@@
Based on the prior model provided in (22), the minimum mean square error (MMSE) estimator for the $(n,k)$-th entry of $\mathbf{X}$ can be expressed as
$\hat{x}_{n,k}=\int{x}_{n,k}p(\beta,\mathbf{Z},\mathbf{X},\boldsymbol{\Gamma},\boldsymbol{\Omega},\mathbf{S}|\mathbf{R})\mathrm{d}\beta\mathrm{d}\mathbf{Z}\mathrm{d}\mathbf{\boldsymbol{\Gamma}}\mathrm{d}\mathbf{\boldsymbol{\Omega}}\mathrm{d}\mathbf{S}\mathrm{d}\mathbf{X},$ (28)
where $p(\beta,\mathbf{Z},\mathbf{X},\boldsymbol{\Gamma},\boldsymbol{\Omega},\mathbf{S}|\mathbf{R})$ denotes the joint posterior probability of $\beta$, $\mathbf{Z}$, $\mathbf{X}$, $\boldsymbol{\Gamma}$, $\boldsymbol{\Omega}$, and $\mathbf{S}$, satisfying
@@
where $p(r_{m,k}|z_{m,k})=\mathcal{CN}(z_{m,k};r_{m,k},\beta^{-1})$; ${r}_{m,k}$ and ${z}_{m,k}$ denote the $(m,k)$-th elements of $\mathbf{R}$ and $\mathbf{Z}$. The dependencies of the random variables in the factorization (29) can be shown by a factor graph as depicted in Fig. 3, where the involved probability distributions are listed in Table I.
Due to the large number of antennas and subcarriers in XL-MIMO systems, solving problem (28) requires evaluating high-dimensional integrals, which is computationally prohibitive. Moreover, traditional SBL-based methods are infeasible in this context as they involve high-dimensional matrix inversions. In addition, the introduction of the common-delay sparsity mechanism and the MRF results in a more complex factor graph representation, making existing message passing algorithms [32, 35, 33, 27] less suitable. Specifically, [32, 35, 33] do not incorporate the common-delay sparsity mechanism or MRF-based variance state modeling, leading to a mismatch between these algorithms and the factor graph representation in Fig. 3. Additionally, [27] lacks the common-delay sparsity mechanism and uses a first-moment matching method to update the prior parameters, which neglects higher-order dependencies among variable nodes. To address these limitations, we propose a novel MMV-VMP algorithm utilizing the variational message passing and belief propagation in the next section.
@@
V Proposed MMV-VMP Algorithm
The MMV-VMP algorithm iteratively updates messages between adjacent nodes and aggregates them at the nodes $x_{n,k}$ to compute the corresponding posterior distributions, thereby circumventing the computationally intensive high-dimensional integrals in (28). In the following sections, we introduce the details of the forward and backward message passing in measurement and prior modules as shown in Fig. 3.
V-A Measurement Module
Measurement module aims to obtain the likelihood estimation of $x_{n,k}$ based on the observation $\mathbf{R}$. Denote the belief of $z_{m,k}$ as $b(z_{m,k})\sim\mathcal{CN}(z_{m,k};\hat{z}_{m,k},\hat{\nu}^{z}_{m,k})$, which is defined in (38) with $\hat{z}_{m,k}$ and $\hat{\nu}^{z}_{m,k}$ indicating the $m$-th elements of $\hat{\mathbf{z}}_{k}$ and $\hat{\boldsymbol{\nu}}_{\mathbf{z}_{k}}$, respectively. In this manner, according to the variation message passing rule, the message from $f_{r_{m,k}}$ to $\beta$ is given by
@@ V-A Measurement Module
Similar to (32), the forward message passing from $f_{r_{m,k}}$ to $z_{m,k}$ can be given by
@@
In this case, utilizing the sum-product rule, the belief of $\mathbf{z}_{k}$ is given by
$b(\mathbf{z}_{k})$ $=\mathcal{CN}(\mathbf{z}_{k};\mathbf{p}_{k},\boldsymbol{\nu}_{\mathbf{p}_{k}})\prod_{m=1}m_{f_{r_{m,k}}\rightarrow z_{m,k}}(z_{m,k})$ (38)
@@
V-B Sparse Prior Module
@@
V-B1 Forward Message Passing
Denote the belief of $x_{n,k}$ as $b(x_{n,k};\hat{x}_{n,k},\hat{\nu}_{\mathbf{x}_{k}})$, defined in (54). Utilizing the variational message passing rule, similar to (32), the message from $f_{x_{n,k}}$ to $\gamma_{n,k}$ is given by
@@ V-B1 Forward Message Passing
Further, according to sum-product rule, the message from $f_{\gamma_{n,k}}$ to $s_{n,k}$ is given by
$m_{f_{\gamma_{n,k}}\rightarrow s_{n,k}}\overset{(a)}{\propto}\int f_{\gamma_{n,k}}m_{f_{x_{n,k}}\rightarrow\gamma_{n,k}}b(\boldsymbol{\alpha}_{k})\mathrm{d}\gamma_{n,k}\mathrm{d}\boldsymbol{\alpha}_{k}$ (43)
@@
Utilizing the message $m_{f_{\gamma_{n,k}}\rightarrow s_{n,k}}(s_{n,k})$, we further derive the message update in the MRF with a 4-connect scheme, where $s_{n,k}^{{\text{l}}}\triangleq s_{n,k-1}$, $s_{n,k}^{\text{r}}\triangleq s_{n,k+1}$, $s_{n,k}^{{\text{t}}}\triangleq s_{n-1,k}$, and $s_{n,k}^{{\text{b}}}\triangleq s_{n,k+1}$ denote the left, right, top and bottom neighbors of $x_{n,k}$. The input message of $s_{n,k}$ from left, right, top and bottom neighbors, denoted as $m_{n,k}^{\mathrm{l}}$, $m_{n,k}^{\mathrm{r}}$, $m_{n,k}^{\mathrm{t}}$, and $m_{n,k}^{\mathrm{b}}$, are Bernoulli distributions. Take $m_{n,k}^{\mathrm{l}}$ as an example, according to the sum-product rule, we have $m_{n,k}^{\mathrm{l}}\propto\lambda_{n,k}^{\mathrm{l}}\delta(1-s_{n,k})+(1-\lambda_{n,k}^{\mathrm{l}})\delta(1+s_{n,k})$, where $\lambda_{n,k}^{\mathrm{l}}$ is given by (45), as shown in the top of next page.
@@
The other three messages can be obtained in a similar way.
V-B2 Backward Message Passing
With the messages of neighbors and $v(s_{n,k})$, the message from $s_{n,k}$ to $f_{\gamma_{n,k}}$ can be given by
@@ V-B2 Backward Message Passing
$\propto\pi^{\mathrm{in}}_{n,k}\delta(1-s_{n,k})+(1-\pi^{\mathrm{in}}_{n,k})\delta(1+s_{n,k})$
@@
According to sum-product rule, the message from $f_{\gamma_{n,k}}$ to $\alpha_{k}^{1}$ is given by
@@
${b}(\alpha^{1}_{k})\propto p(\alpha_{k}^{1})\prod_{n}m_{f_{\gamma_{n,k}}\rightarrow\alpha_{k}^{1}}(\alpha_{k}^{1})\propto\mathrm{Ga}(\alpha_{k}^{1};\hat{a}_{k}^{1},\hat{b}_{k}^{1}),$ (49)
where $\hat{a}_{k}^{1}=a_{1}+N$ and $\hat{b}_{k}^{1}=b_{1}+\sum_{n=1}^{N}\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}}$. Similarly, we have ${b}(\alpha^{2}_{k})\propto\mathrm{Ga}(\alpha_{k}^{2};\hat{a}_{k}^{2},\hat{b}_{k}^{2})$ with $\hat{a}_{k}^{2}=a_{2}+N$ and $\hat{b}_{k}^{2}=b_{2}+\sum_{n=1}^{N}\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}}$. According to the sum-product rule, the message from $f_{\gamma_{n,k}}$ to $\gamma_{n,k}$ is given by
@@
As a result, the message from $f_{x_{n,k}}$ to $x_{n,k}$ is denoted as
$m_{f_{x_{n,k}}\rightarrow x_{n,k}}(x_{n,k})$ $\propto\exp\left\{\int b(\gamma_{n,k})\ln f_{x_{n,k}}\mathrm{d}\gamma_{n,k}\right\}$ (52)
@@
$\hat{\gamma}_{n,j}$ $=\int\gamma_{n,k}b(\gamma_{n,k})\mathrm{d}\gamma_{n,k}=\pi^{\mathrm{in}}_{n,k}\frac{\hat{a}_{k}^{1}+1}{\hat{b}_{k}^{1}+\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}}}$ (53)
@@
Algorithm 1 Proposed UAMP-SBL-MRF algorithm
@@
2:Initialize: $\hat{\nu}_{\mathbf{x}_{k}}^{(0)}=1$, $\hat{\mathbf{x}}=\mathbf{0}$, $\hat{{\gamma}}_{n,k}={1},\hat{\beta}=1$, and $\boldsymbol{\mu}=\mathbf{0}$.
@@
11:     Update the messages $m_{f_{x_{n,k}}\rightarrow\gamma_{n,k}}$ according to (43);
12:     Update the messages $m_{n}^{\mathrm{l}}$, $m_{n}^{\mathrm{r}}$, $m_{n}^{\mathrm{t}}$, and $m_{n}^{\mathrm{b}}$;
13:     Update the messages $m_{s_{n}\rightarrow f_{\gamma_{n}}}(s_{n})$ according to (47);
@@
15:     Update $\hat{\gamma}_{n}$ according to (53);
@@
The proposed MMV-VMP algorithm is summarized in Algorithm 1 and it can be terminated when it reached a maximum number of iteration or the difference between the estimates of two consecutive iterations is less than $10^{-5}$. In the following, we provide the computational complexity analysis for the proposed MMV-VMP algorithm. Examining the steps of Algorithm 1, it is evident that there is no matrix inversion involved. Thus, the most computationally intensive parts only involve matrix-vector products in lines 2 and 6, i.e., $\mathcal{O}(MN)$ per iteration. Consequently, the total complexity of the MMV-VMP algorithm is $\mathcal{O}(TKMN)$, where $T$ denotes the number of iterations.

### v3

@@ IV-B Column-Wise Hierarchical Sparse Prior Model
According to Lemma 1 and Theorem 1, the angular-delay domain channel exhibits prominent block sparsity and a column-wise clustered structure. However, existing sparse prior models fail to capture these characteristics fully [40, 39, 42, 41]. For example, the two-layer hierarchical prior models in [40, 39] assign independent priors to the variance precision parameters of the channel coefficients, which limits their ability to model the correlation across coefficients. Although [42, 41] leverage MRF to promote global block sparsity, they still overlook local structural dependencies, especially those induced by column-wise clustering. As a result, these models are not well suited for the SnS wideband XL-MIMO channel estimation problem considered in this work.
To this end, we propose a novel column-wise hierarchical sparse prior model, which incorporates both precision parameter sharing mechanism and MRF structure to promote structured sparsity. Specifically, the column-wise hierarchical prior is defined as
$p(\mathbf{X},\boldsymbol{\Gamma},\boldsymbol{\Omega},\mathbf{S})=p(\mathbf{X}|\boldsymbol{\Gamma})p(\boldsymbol{\Gamma}|\boldsymbol{\Omega},\mathbf{S})p(\boldsymbol{\Omega})p(\mathbf{S}).$ (31)
We next detail the distributions associated with each layer.
1) Variance-Driven Sparsity: The first layer adopts a complex Gaussian prior with zero mean and variance $[\boldsymbol{\Gamma}]_{n,k}=\gamma_{n,k}^{-1}$, i.e.,
$p(\mathbf{X}|\boldsymbol{\Gamma})=\prod_{n=1}^{N}\prod_{k=1}^{K}\mathcal{CN}(x_{n,k};0,\gamma_{n,k}^{-1}),$ (32)
where $\gamma_{n,k}$ denotes the variance precision parameter, with $\gamma_{n,k}^{-1}$ representing the variance of $x_{n,k}$. This precision-based modeling naturally induces sparsity: as the variance $\gamma_{n,k}^{-1}$ approaches zero, the corresponding $x_{n,k}$ is effectively pushed toward zero, thereby promoting a sparse solution.
2) Precision Parameters Sharing Mechanisms: Due to the finite angular and delay resolutions, the angular-delay domain channel does not exhibit strict sparsity, as the coefficients near the boundaries of the sparse patterns often retain non-negligible energy. Enforcing hard zeros on these components may lead to degraded estimation performance. To mitigate this issue, in the second layer, we adopt a dual-precision parameter strategy, wherein each variance $\gamma_{n,k}$ is associated with two distinct precision parameters corresponding to its active and inactive states. In addition, to capture the local common-delay clustered sparsity, we further introduce the precision parameters sharing mechanism. Specifically, $\gamma_{n,k}$ is assigned a conditionally Bernoulli-Gaussian distribution given by
@@
$p(\boldsymbol{\Omega})=\prod_{i=1}^{2}\prod_{k=1}^{K}p(\alpha_{k}^{i})=\prod_{i=1}^{2}\prod_{k=1}^{K}\mathrm{Ga}(\alpha_{k}^{i};a_{i},b_{i}).$ (34)
As described in (33) and (34), this common-delay sparsity mechanism ensures that, for a fixed $k$, all $x_{n,k}$ sharing the same variance state are governed by a common variance precision parameter. This coupling enforces a consistent sparsity structure among components associated with the same delay.
3) MRF-Based Variance State Modeling: In the third layer, to capture the global block sparsity, the variance state variable $\mathbf{S}$ is modeled as a MRF
@@
where $\mathcal{D}{n,k}$ denotes the set of neighboring nodes of $s_{n,k}$; $u(s_{n,k},s_{n^{\prime},k^{\prime}})=\exp(\varpi s_{n,k}s_{n^{\prime},k^{\prime}})$ and $v(s_{n,k})=\exp(-\eta s_{n,k})$ represent the pairwise potential and self-potential functions, respectively, with $\varpi$ and $\eta$ being the model parameters associated with $p(\mathbf{S})$. By leveraging the formulation in (35), the prior encourages block patterns in the support structure and suppresses isolated coefficients that deviate from their neighboring states.
V Proposed MMV-HMP Algorithm
Based on the proposed column-wise hierarchical sparse prior model, this section first formulates the sparse signal recovery problem as a MMV-based Bayesian inference task. Then, to effectively perform the inference, we propose an MMV-HMP algorithm.
V-A Bayesian Inference
@@
Based on the prior model provided in (31), the maximum a posterior (MAP) estimator for the $(n,k)$-th entry of $\mathbf{X}$ can be expressed as
$\hat{x}_{n,k}=\int{x}_{n,k}p(x_{n,k}|\mathbf{R})\mathrm{d}x_{n,k},$ (37)
where the posterior distribution $p(x_{n,k}|\mathbf{R})$ is given by
$\int p(\beta,\mathbf{Z},\mathbf{X},\boldsymbol{\Gamma},\boldsymbol{\Omega},\mathbf{S}|\mathbf{R})\,\mathrm{d}\beta\,\mathrm{d}\mathbf{Z}\,\mathrm{d}\mathbf{\boldsymbol{\Gamma}}\,\mathrm{d}\mathbf{\boldsymbol{\Omega}}\,\mathrm{d}\mathbf{S}\,\mathrm{d}\mathbf{X}_{\backslash n,k}$ (38)
with $p(\beta,\mathbf{Z},\mathbf{X},\boldsymbol{\Gamma},\boldsymbol{\Omega},\mathbf{S}|\mathbf{R})$ indicating the joint posterior probability and satisfying
@@
where $p(r_{m,k}|z_{m,k})=\mathcal{CN}(z_{m,k};r_{m,k},\beta^{-1})$; ${r}_{m,k}$ and ${z}_{m,k}$ denote the $(m,k)$-th elements of $\mathbf{R}$ and $\mathbf{Z}$.
Due to the large number of antennas and subcarriers in XL-MIMO systems, solving problem (37) requires evaluating high-dimensional integrals, which is computationally prohibitive. Moreover, traditional SBL-based methods are infeasible in this context as they involve high-dimensional matrix inversions. Consequently, this highlights the pressing need for novel channel estimation techniques capable of addressing the challenges posed by the SnS dual-wideband channel estimation in XL-MIMO systems. Recently, message passing-based techniques have been widely adopted for solving MAP estimation problems, owing to their computational efficiency. Motivated by these advances, we propose a MMV-HMP algorithm to efficiently solve problem (37).
V-B Factor Graph Representation
The dependencies among the random variables in the factorization (39) are illustrated using a factor graph, as shown in Fig. 3, with the corresponding probability distributions summarized in Table I. It is evident that the factor graph in Fig. 3 is significantly more intricate than those considered in [39, 43, 40, 42, 41], primarily due to the incorporation of precision parameter sharing mechanism and MRF structure. Specifically, the works in [39, 43, 40] adopt independent priors on the variances of coefficients, thereby promoting element-wise sparsity. However, such models fail to capture the angular or delay-domain correlations among the coefficients. Furthermore, the model in [42] does not include a precision parameter sharing mechanism, and thus overlooks higher-order dependencies among variable nodes. Due to these structural distinctions, existing message passing algorithms developed in [39, 43, 40, 42, 41] are not directly applicable to our setting. Consequently, the message update equations at the variable nodes must be carefully reformulated to accommodate the enhanced complexity of our factor graph.
@@ V-B Factor Graph Representation
Moreover, in the proposed sparse prior model, the latent precision parameters $\beta$ and $\alpha^{i}_{k}$ follow non-Gaussian distributions. Meanwhile, the shared precision parameter simultaneously influences the message updates in both the first and third layers, resulting in inter-layer dependencies, which significantly increase the complexity of inference and render the exact computation of the node beliefs computationally intractable under a pure belief propagation (BP) framework.
To overcome these challenges, we propose a novel MMV-HMP algorithm jointly leveraging VMP and sum-produce algorithm (SPA) rules. Specifically, tailored to the intractable structure of the factor graph, we reformulate the message update equations along the edges, selectively applying sum-product or variational message passing rules to different edges to enable more flexible and efficient message updates.
The MMV-HMP algorithm iteratively updates messages between adjacent nodes and aggregates them at the nodes $x_{n,k}$ to compute the corresponding posterior distributions, thereby circumventing the computationally intensive high-dimensional integrals in (37). In the following sections, we introduce the details of the forward and backward message passing in measurement and prior modules as shown in Fig. 3.
V-C Measurement Module
Measurement module aims to obtain the likelihood estimation of $x_{n,k}$ based on the observation $\mathbf{R}$. Denote the belief of $z_{m,k}$ as $b(z_{m,k})\sim\mathcal{CN}(z_{m,k};\hat{z}_{m,k},\hat{\nu}^{z}_{m,k})$, which is defined in (48) with $\hat{z}_{m,k}$ and $\hat{\nu}^{z}_{m,k}$ indicating the $m$-th elements of $\hat{\mathbf{z}}_{k}$ and $\hat{\boldsymbol{\nu}}_{\mathbf{z}_{k}}$, respectively. In this manner, according to the VMP rule [44, 45], the message from $f_{r_{m,k}}$ to $\beta$ is given by
@@ V-C Measurement Module
Similar to (42), utilizing VMP rule, the forward message passing from $f_{r_{m,k}}$ to $z_{m,k}$ can be given by
@@
In this case, utilizing SPA rule, the belief of $\mathbf{z}_{k}$ is given by
$b(\mathbf{z}_{k})$ $=\mathcal{CN}(\mathbf{z}_{k};\mathbf{p}_{k},\boldsymbol{\nu}_{\mathbf{p}_{k}})\prod_{m=1}m_{f_{r_{m,k}}\rightarrow z_{m,k}}$ (48)
@@
V-D Sparse Prior Module
@@
V-D1 Forward Message Passing
Denote the belief of $x_{n,k}$ as $b(x_{n,k};\hat{x}_{n,k},\hat{\nu}_{\mathbf{x}_{k}})$, defined in (63). Utilizing VMP rule, similar to (42), the message from $f_{x_{n,k}}$ to $\gamma_{n,k}$ is given by
@@ V-D1 Forward Message Passing
Furthermore, according to SPA rule, the message from $f_{\gamma_{n,k}}$ to $s_{n,k}$ is given by
$m_{f_{\gamma_{n,k}}\rightarrow s_{n,k}}(s_{n,k})\overset{(a)}{\propto}\int f_{\gamma_{n,k}}m_{f_{x_{n,k}}\rightarrow\gamma_{n,k}}b(\boldsymbol{\alpha}_{k})\mathrm{d}\gamma_{n,k}\mathrm{d}\boldsymbol{\alpha}_{k}$ (53)
@@
Utilizing the message $m_{f_{\gamma_{n,k}}\rightarrow s_{n,k}}(s_{n,k})$, we further derive the message update in the MRF with a 4-connect scheme, where $s_{n,k}^{{\text{l}}}\triangleq s_{n,k-1}$, $s_{n,k}^{\text{r}}\triangleq s_{n,k+1}$, $s_{n,k}^{{\text{t}}}\triangleq s_{n-1,k}$, and $s_{n,k}^{{\text{b}}}\triangleq s_{n,k+1}$ denote the left, right, top and bottom neighbors of $x_{n,k}$. The input message of $s_{n,k}$ from left, right, top and bottom neighbors, denoted as $m_{n,k}^{\mathrm{l}}$, $m_{n,k}^{\mathrm{r}}$, $m_{n,k}^{\mathrm{t}}$, and $m_{n,k}^{\mathrm{b}}$, are Bernoulli distributions. Take $m_{n,k}^{\mathrm{l}}$ as an example, according to SPA rule, we have $m_{n,k}^{\mathrm{l}}\propto\lambda_{n,k}^{\mathrm{l}}\delta(1-s_{n,k})+(1-\lambda_{n,k}^{\mathrm{l}})\delta(1+s_{n,k})$, where $\lambda_{n,k}^{\mathrm{l}}$ is given by (54), as shown in the top of this page.
@@
The other messages can be obtained in a similar way.
V-D2 Backward Message Passing
With the messages of neighbors and $v(s_{n,k})$ and SPA rule, the message from $s_{n,k}$ to $f_{\gamma_{n,k}}$ can be given by
@@ V-D2 Backward Message Passing
$\propto$ $\pi^{\mathrm{in}}_{n,k}\delta(1-s_{n,k})+(1-\pi^{\mathrm{in}}_{n,k})\delta(1+s_{n,k}),$
@@
According to SPA rule, the message from $f_{\gamma_{n,k}}$ to $\alpha_{k}^{1}$ is given by
@@
${b}(\alpha^{1}_{k})\propto p(\alpha_{k}^{1})\prod_{n}m_{f_{\gamma_{n,k}}\rightarrow\alpha_{k}^{1}}\propto\mathrm{Ga}(\alpha_{k}^{1};\hat{a}_{k}^{1},\hat{b}_{k}^{1}),$ (58)
where $\hat{a}_{k}^{1}=a_{1}+N$ and $\hat{b}_{k}^{1}=b_{1}+\sum_{n=1}^{N}\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}}$. Similarly, we have ${b}(\alpha^{2}_{k})\propto\mathrm{Ga}(\alpha_{k}^{2};\hat{a}_{k}^{2},\hat{b}_{k}^{2})$ with $\hat{a}_{k}^{2}=a_{2}+N$ and $\hat{b}_{k}^{2}=b_{2}+\sum_{n=1}^{N}\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}}$. According to SPA rule, the message from $f_{\gamma_{n,k}}$ to $\gamma_{n,k}$ is given by
@@
As a result, utilizing VMP rule, the message from $f_{x_{n,k}}$ to $x_{n,k}$ is denoted as
$m_{f_{x_{n,k}}\rightarrow x_{n,k}}$ $\propto\exp\left\{\int b(\gamma_{n,k})\ln f_{x_{n,k}}\mathrm{d}\gamma_{n,k}\right\}$ (61)
@@
$\hat{\gamma}_{n,k}$ $=\int\gamma_{n,k}b(\gamma_{n,k})\mathrm{d}\gamma_{n,k}=\pi^{\mathrm{in}}_{n,k}\frac{\hat{a}_{k}^{1}+1}{\hat{b}_{k}^{1}+\left|\hat{x}_{n,k}\right|^{2}+\hat{\nu}_{\mathbf{x}_{k}}}$ (62)
@@
Algorithm 1 Proposed MMV-HMP algorithm
@@
2:Initialize: $\hat{\nu}_{\mathbf{x}_{k}}^{(0)}=1$, $\hat{\mathbf{x}}_{k}=\mathbf{0}$, $\hat{{\gamma}}_{n,k}={1},\hat{\beta}=1$, and $\boldsymbol{\mu}_{k}=\mathbf{0}$.
@@
11:     Update the messages $m_{f_{x_{n,k}}\rightarrow s_{n,k}}$ according to (53);
12:     Update the messages $m_{n,k}^{\mathrm{l}}$, $m_{n,k}^{\mathrm{r}}$, $m_{n,k}^{\mathrm{t}}$, and $m_{n,k}^{\mathrm{b}}$;
13:     Update the messages $m_{s_{n,k}\rightarrow f_{\gamma_{n,k}}}$ according to (56);
@@
15:     Update $\hat{\gamma}_{n,k}$ according to (62);
@@
The proposed MMV-HMP algorithm is summarized in Algorithm 1 and it can be terminated when it reached a maximum number of iteration or the difference between the estimates of two consecutive iterations is less than $10^{-5}$. In the following, we provide the computational complexity analysis for the proposed MMV-HMP algorithm. Examining the steps of Algorithm 1, it is evident that there is no matrix inversion involved. Thus, the most computationally intensive parts only involve matrix-vector products in lines 2 and 6, i.e., $\mathcal{O}(MN)$ per iteration. Consequently, the total complexity of the MMV-HMP algorithm is $\mathcal{O}(TKMN)$, where $T$ denotes the number of iterations.
VI Simulation Results